# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v2

@@ Abstract
The NP-hard Dynamic Flexible Job-Shop Scheduling (DFJSP) problem involves real-time events and complex routing. While traditional rules are efficient but rigid, deep learning is opaque and requires feature engineering. Large Language Models (LLMs) promise adaptive reasoning without this engineering overhead, yet we find their direct application is suboptimal. Baseline LLMs suffer from three key pitfalls: the long-context paradox, where crucial data is underutilized; an underutilization of expert heuristics; and myopic decision-making. To address this, we propose ReflecSched, a framework that empowers the LLM beyond a direct scheduler by equipping it with a strategic analysis capability. ReflecSched tasks the LLM to analyze heuristic-driven simulations across multiple planning horizons and distill them into a concise, natural-language summary termed “Strategic Experience”. This summary is then integrated into the prompt of a final decision-making module, guiding it to produce non-myopic actions. Experiments demonstrate ReflecSched achieves superior performance, with its best variants attaining an average RPD of 6.04% and rank of 3.18, significantly outperforming strong traditional and learning-based methods. It also statistically and decisively surpasses direct LLM baselines, securing a 71.35% Win Rate while being, on average, 15.1% more token-efficient on Normal-scale problems. Ablation studies attribute this performance to a robust reflection mechanism that leverages high-quality, contrastive experience. This mechanism mitigates key LLM pitfalls like myopic greed, enabling ReflecSched to outperform all evaluated heuristics. Ultimately, the framework’s performance is statistically on par with an oracle-like strategy, showcasing its effectiveness and robustness.
@@ 1 Introduction
Dynamic Flexible Job-Shop Scheduling (DFJSP) is a long-standing challenge in operations research and a key component of modern smart manufacturing [li2025learning]. As an NP-hard problem, it involves the complex task of continuously allocating operations to machines in a real-time environment characterized by stochastic events such as new job arrivals or machine breakdowns. The ability to efficiently schedule these operations critically impacts the agility, resilience, and profitability of production systems. For decades, the field has been dominated by methods ranging from classical heuristics to meta-heuristics [cao2023inverse]. While foundational, these approaches often rely on handcrafted rules that can struggle to generalize across diverse and unforeseen dynamic scenarios [ferreira2022effective].
In recent years, deep learning (DL), particularly reinforcement learning (RL), has emerged as a promising paradigm for this problem [xu2025learn]. By training agents on simulated data, DL-based methods can learn sophisticated, state-aware scheduling policies. However, this capability comes at a significant cost. A key limitation lies in the intricate process of state and action space engineering: translating the rich state of a dynamic workshop into a fixed-size numerical vector that a neural network can process is a complex task requiring dual expertise in scheduling and machine learning [zhang2020learning]. Furthermore, the resulting models are often “black boxes” with opaque decision-making logic that is difficult for engineers to interpret or trust in high-stakes environments [ngwu2025reinforcement].
The advent of Large Language Models (LLMs) offers a different approach, potentially circumventing these challenges by leveraging natural language as a more accessible interface. An LLM-based scheduler could, in principle, reason directly over a textual description of the factory state, reducing the need for complex numerical encoding [abgaryan2024llms]. This suggests the possibility of developing and adapting scheduling strategies with greater flexibility.
@@ 2 Related Work
Heuristic and Metaheuristic Approaches: Heuristic-based methods are a foundational approach for the Dynamic Job-Shop Scheduling Problem (DJSP) due to their computational tractability. Research has advanced from simple dispatching rules to complex metaheuristics designed to handle diverse operational constraints. For instance, some works have applied GRASP algorithms to manage a wide array of dynamic events [baykasouglu2020greedy], while others have used hybrid Particle Swarm Optimization for joint production and transportation scheduling [ren2022joint]. Further research has focused on specific uncertainties, such as variable processing times, using techniques like Artificial Bee Colony algorithms [shahgholi2019heuristic]. A key limitation of these metaheuristics, however, is their reliance on a pre-defined search logic that is inherently less adaptive and struggles to generalize to new, unseen problem scenarios.
@@
Deep Reinforcement Learning Approaches: To address the generalization limitations of heuristics, Deep Reinforcement Learning (DRL) has emerged as an alternative that learns adaptive policies, typically by formulating the DJSP as a Markov Decision Process. Architectural innovations have been central to this area. For example, Pointer Networks have been utilized for end-to-end policy learning that avoids reliance on predefined rules [yang2024deep]. To enhance scalability, a multi-agent perspective has been adopted through decentralized frameworks [liu2023deep]. Other approaches decompose the decision-making process using hierarchical structures [luo2021real]. Nevertheless, DRL presents its own significant challenges: performance is highly sensitive to the complex engineering of state and reward functions; the resulting policies are often opaque “black boxes” hindering trust; and they demand accurate simulation environments for training, which are often unavailable for complex real-world systems.
@@
Large Language Models Paradigms: Seeking to address the challenges of both handcrafted logic and the black-box nature of DRL, LLMs have recently introduced a new paradigm for scheduling by leveraging their vast pre-existing knowledge and reasoning capabilities. One direction explores LLMs as end-to-end solvers by fine-tuning them on large-scale datasets of problems paired with their solutions [abgaryan2025starjob, abgaryan2024llms]. Another approach employs LLMs as automated problem modelers that translate natural language into executable code for traditional solvers [amarasinghe2023ai, ahmaditeshnizi2024optimus]. A third direction positions the LLM as a hyper-heuristic generator that iteratively reflects upon and improves heuristic rule-code [huang2024automatic]. Despite their potential, early works reveal critical challenges: the risk of generating infeasible solutions (“hallucinations”), high computational costs, and a primary focus on static problems, leaving their application to complex, dynamic scheduling environments largely unexplored.
@@
Hierarchical Memory and Reasoning Paradigms: A significant body of work explores hierarchical paradigms to structure and access information beyond the fixed context window. A prominent direction involves creating hierarchical memory systems, such as MemGPT’s OS-inspired information paging [packer2024memgptllmsoperatingsystems] or the semantic-level memory structures in MemoryBank and H-MEM that enable targeted retrieval [zhong2023memorybankenhancinglargelanguage, sun2025hierarchicalmemoryhighefficiencylongterm]. A related approach, seen in HIAGENT, employs hierarchy for task decomposition to manage an agent’s working memory during long-horizon tasks [hu2024hiagenthierarchicalworkingmemory]. Diverging from these methods, which primarily focus on organizing and retrieving an agent’s past, our proposed ReflecSched framework introduces a hierarchical reflection mechanism that functions as a dynamic, planning-time operator. Instead of retrieving historical data, it prospectively explores the future by leveraging multi-level, heuristic-driven simulations to analyze potential decision pathways. The hierarchy in ReflecSched is thus one of planning abstraction distilling numerical simulations into a transient, textual Strategic Experience to guide the next action, rather than creating a persistent memory store.
@@ 3.1 Dynamic Flexible Job-Shop Scheduling
DFJSP is the problem of assigning operations from multiple jobs to a set of machines, subject to real-time stochastic events. It is characterized by two features: flexibility, which allows each operation to be processed on a subset of candidate machines, and dynamism, which refers to the occurrence of real-time stochastic events such as new job arrivals or machine breakdowns. The objective is to generate a valid schedule that minimizes the makespan, defined as the maximum completion time among all jobs [cao2024novel]. Figure 1 illustrates an example of such a schedule, visualized as a Gantt chart. A valid schedule must satisfy both precedence constraints between operations of the same job and the resource constraints of the machines.
In this work, we model a rich set of dynamic events to capture the stochastic nature of modern manufacturing environments. These include new job arrivals, machine breakdowns and subsequent repairs, job cancellations, and the introduction of high-priority emergency jobs. This set of events is representative of the primary disruptions studied in the DFJSP literature [baykasoglu2020greedy, ngwu2025reinforcement]. More importantly, the core architecture of ReflecSched is designed to be event-agnostic and generalizable. Its strategic Hierarchical Reflection module is invoked by an event-driven trigger, which responds to any significant state change rather than to a specific event type. Consequently, the framework can be readily extended to handle other stochastic events, such as variable processing times [shahgholi2019heuristic] or rush order arrivals [ren2024dynamic], without requiring modifications to its fundamental logic.
@@ 3.2 Mathematical Formulation of the DFJSP
Table 1: Notation for the DFJSP Mathematical Formulation
Symbol Description
$\mathcal{J}$ Set of active jobs currently in the system, $\{1,...,n\}$.
$\mathcal{M}$ Set of available machines, $\{1,...,m\}$.
$\mathcal{O}_{i}$ Set of operations for job $J_{i}\in\mathcal{J}$.
$\mathcal{O}$ Set of all operations to be scheduled.
$M_{ij}$ Set of candidate machines for operation $O_{ij}$.
$n_{i}$ Number of operations for job $J_{i}$.
$p_{ijk}$ Processing time of $O_{ij}$ on machine $k\in M_{ij}$.
$a_{i}$ Arrival time of job $J_{i}$.
$r_{k}(t)$ Earliest time machine $k$ is free.
$M$ A sufficiently large positive number (a “Big-M” constant).
$s_{ij}$ Start time of operation $O_{ij}$.
$c_{ij}$ Completion time of operation $O_{ij}$.
$x_{ijk}$ Binary: 1 if $O_{ij}$ is assigned to machine $k$.
$C_{max}$ The makespan, defined as $\max_{i\in\mathcal{J}}\{c_{i,n_{i}}\}$.
@@ 3.2.1 Objective Function and Constraints
$s_{i^{\prime}j^{\prime}}$ $\geq c_{ij}-M\cdot(3-x_{ijk}-x_{i^{\prime}j^{\prime}k}-y_{ij,i^{\prime}j^{\prime}k})$ (7)
$s_{ij}$ $\geq c_{i^{\prime}j^{\prime}}-M\cdot(2+y_{ij,i^{\prime}j^{\prime}k}-x_{ijk}-x_{i^{\prime}j^{\prime}k})$ (8)
@@ 3.2.2 Formalization of Dynamic Events
The dynamic nature of the DFJSP is addressed by updating this formulation in response to stochastic events. Let the state of the system at time $t$ be defined by the tuple $\Sigma_{t}=(\mathcal{J}_{t},\{a_{i}\},\{r_{k}(t)\},\mathcal{O}_{int})$. A dynamic event at $t_{event}$ triggers a state transition from $\Sigma_{t_{event}}$ to $\Sigma_{t_{event}^{+}}$, where $t_{event}^{+}$ denotes the time immediately after the event. The updates for each event type are formalized below [baykasouglu2020greedy].
Job Arrival: At time $t_{event}=t_{arr}$, a new job $J_{new}$ with its own set of operations $\mathcal{O}_{new}$ and arrival time $a_{new}=t_{arr}$ is introduced into the system. The state is updated as follows:
$\mathcal{J}_{t_{arr}^{+}}$ $\leftarrow\mathcal{J}_{t_{arr}}\cup\{J_{new}\}$ (9)
$\mathcal{O}_{t_{arr}^{+}}$ $\leftarrow\mathcal{O}_{t_{arr}}\cup\mathcal{O}_{new}$ (10)
The emergency status of a job is a characteristic handled by the scheduling policy (e.g., via prioritization) rather than being a direct modification to the model’s constraints.
Job Cancellation: At time $t_{event}=t_{cancel}$, a job $J_{cancel}$ is removed from the system. This action removes the job from the set of active jobs $\mathcal{J}$ and all of its corresponding uncompleted operations from the set of operations to be scheduled $\mathcal{O}$. The historical record of any completed operations of $J_{cancel}$ is preserved, but they are no longer relevant for future scheduling decisions.
@@
$\mathcal{J}_{t_{cancel}^{+}}$ $\leftarrow\mathcal{J}_{t_{cancel}}\setminus\{J_{cancel}\}$ (11)
$\mathcal{O}_{t_{cancel}^{+}}$ $\leftarrow\mathcal{O}_{t_{cancel}}\setminus\mathcal{O}_{cancel}$ (12)
If an operation from $\mathcal{O}_{cancel}$ was in process on a machine $k$ at $t_{cancel}$, it is preempted. The machine is immediately freed, and its availability time $r_{k}(t_{cancel}^{+})$ is updated to $t_{cancel}$.
Machine Breakdown: At time $t_{event}=t_{break}$, a machine $k\in\mathcal{M}$ fails and becomes unavailable. The immediate consequence is that the machine’s earliest availability time is postponed until its projected repair time, $t_{repair}$. This is formalized as:
$r_{k}(t_{break}^{+})=t_{repair}$ (13)
This update affects all subsequent scheduling decisions, as no new operations can be assigned to machine $k$ until $t_{repair}$.
If an operation $O_{ij}$ was being processed on machine $k$ at this time (i.e., $s_{ij}<t_{break}<c_{ij}$), it is immediately interrupted. The operation is formally moved to a set of interrupted operations, $\mathcal{O}_{int}$, to signify its special status. It retains its assignment to machine $k$ and awaits the machine’s repair to resume processing.
Machine Repair: At time $t_{event}=t_{repair}$, machine $k$ becomes operational again. Any operation $O_{ij}\in\mathcal{O}_{int}$ that was interrupted on this machine can now resume. Upon repair, $O_{ij}$ is removed from the set $\mathcal{O}_{int}$.
@@
$c^{\prime}_{ij}=t_{repair}+p_{ijk}-(t_{break}-s_{ij})$ (14)
where $t_{break}-s_{ij}$ represents the processing work completed before the interruption. This new completion time $c^{\prime}_{ij}$ subsequently serves as the basis for re-evaluating the precedence constraints for all downstream operations of job $J_{i}$. The machine $k$ is also now available for scheduling other operations from time $t_{repair}$ onwards.

### v3

@@ Abstract
The NP-hard Dynamic Flexible Job-Shop Scheduling (DFJSP) problem involves real-time events and complex routing. While traditional rules are efficient but rigid, deep learning is opaque and requires feature engineering. Large Language Models (LLMs) promise adaptive reasoning without this engineering overhead, yet we find their direct application is suboptimal. Baseline LLMs suffer from three key pitfalls: the long-context paradox, where crucial data is underutilized; an underutilization of expert heuristics; and myopic decision-making. To address this, we propose ReflecSched, a framework that empowers the LLM beyond a direct scheduler by equipping it with a strategic analysis capability. ReflecSched tasks the LLM to analyze heuristic-driven simulations across multiple planning horizons and distill them into a concise, natural-language summary termed “Strategic Experience”. This summary is then integrated into the prompt of a final decision-making module, guiding it to produce non-myopic actions. Experiments demonstrate ReflecSched achieves superior performance, with its best variants attaining an average RPD of 6.09% and rank of 4.39 on GEN-Bench, significantly outperforming strong traditional and learning-based methods including HMPSAC and IDDQN. Extensive cross-benchmark evaluations on MK-Bench and the semiconductor-focused JMS-Bench further confirm the framework’s robustness, yielding minimum average RPDs of 6.83% and 6.18%, respectively. It also statistically and decisively surpasses direct LLM baselines, securing a 71.35% Win Rate while being, on average, 15.1% more token-efficient on Normal-scale problems. Furthermore, cumulative runtime analysis reveals that ReflecSched’s zero-shot nature eliminates the training bottleneck, providing a decisive efficiency advantage in high-variability manufacturing environments. Ablation studies attribute this performance to a robust reflection mechanism that leverages high-quality, contrastive experience. This mechanism mitigates key LLM pitfalls like myopic greed, enabling ReflecSched to outperform all evaluated heuristics. Ultimately, the framework’s performance is statistically on par with an oracle-like strategy, showcasing its effectiveness and robustness.
@@ 1 Introduction
Dynamic Flexible Job-Shop Scheduling (DFJSP) is a long-standing challenge in operations research and a key component of modern smart manufacturing [28]. As an NP-hard problem, it involves the complex task of continuously allocating operations to machines in a real-time environment characterized by stochastic events such as new job arrivals or machine breakdowns. The ability to efficiently schedule these operations critically impacts the agility, resilience, and profitability of production systems. For decades, the field has been dominated by methods ranging from classical heuristics to meta-heuristics [11]. While foundational, these approaches often rely on handcrafted rules that can struggle to generalize across diverse and unforeseen dynamic scenarios [17].
In recent years, deep learning (DL), particularly reinforcement learning (RL), has emerged as a promising paradigm for this problem [59]. By training agents on simulated data, DL-based methods can learn sophisticated, state-aware scheduling policies. However, this capability comes at a significant cost. A key limitation lies in the intricate process of state and action space engineering: translating the rich state of a dynamic workshop into a fixed-size numerical vector that a neural network can process is a complex task requiring dual expertise in scheduling and machine learning [66]. Furthermore, the resulting models are often “black boxes” with opaque decision-making logic that is difficult for engineers to interpret or trust in high-stakes environments [40].
The advent of Large Language Models (LLMs) offers a different approach, potentially circumventing these challenges by leveraging natural language as a more accessible interface. An LLM-based scheduler could, in principle, reason directly over a textual description of the factory state, reducing the need for complex numerical encoding [2]. This suggests the possibility of developing and adapting scheduling strategies with greater flexibility.
@@
  • 3.
    Multi-tiered benchmark suites and open-source ecosystem. We introduce a comprehensive benchmarking ecosystem comprising GEN-Bench, MK-Bench, and JMS-Bench. These suites span from general-purpose scenarios to standardized flexible job-shop instances and high-fidelity semiconductor manufacturing environments. By open-sourcing these benchmarks alongside the ReflecSched framework, we provide a rigorous, multi-dimensional evaluation protocol to facilitate future research in LLM-based industrial planning.
@@ 2 Related Work
Heuristic and Metaheuristic Approaches: Heuristic-based methods are a foundational approach for the Dynamic Job-Shop Scheduling Problem (DJSP) due to their computational tractability. Research has advanced from simple dispatching rules to complex metaheuristics designed to handle diverse operational constraints. For instance, some works have applied greedy randomized adaptive search algorithms to manage a wide array of dynamic events [7], while others have used hybrid Particle Swarm Optimization for joint production and transportation scheduling [44]. Further research has focused on specific uncertainties, such as variable processing times, using techniques like Artificial Bee Colony algorithms [46]. To align with the human-centric paradigm of Industry 5.0, recent research has formalized the HDDFJSP by integrating worker-related dynamic factors into a multi-objective optimization framework, leveraging a Q-learning-enhanced differential evolution algorithm to balance production efficiency with operator well-being [29]. A key limitation of these metaheuristics, however, is their reliance on a pre-defined search logic that is inherently less adaptive and struggles to generalize to new, unseen problem scenarios [13].
@@
Deep Reinforcement Learning Approaches: To address the generalization limitations of heuristics, Deep Reinforcement Learning (DRL) has emerged as an alternative that learns adaptive policies, typically by formulating the DJSP as a Markov Decision Process. Architectural innovations have been central to this area. For example, Pointer Networks have been utilized for end-to-end policy learning that avoids reliance on predefined rules [61]. To enhance scalability, a multi-agent perspective has been adopted through decentralized frameworks [34]. Other approaches decompose the decision-making process using hierarchical structures [37]. Recently, to enhance the adaptability of dynamic job-shop scheduling, Yu et al. introduced a DRL-based framework that reformulates scheduling states as multi-channel images, leveraging a Spatial Pyramid Pooling Fast module to achieve scale-independent feature extraction and a region-based dense reward function to facilitate fine-grained policy optimization [64]. Nevertheless, DRL presents its own significant challenges: performance is highly sensitive to the complex engineering of state and reward functions; the resulting policies are often opaque “black boxes” hindering trust; and they demand accurate simulation environments for training, which are often unavailable for complex real-world systems.
@@
Large Language Models Paradigms: Seeking to address the challenges of both handcrafted logic and the black-box nature of DRL, LLMs have recently introduced a new paradigm for scheduling by leveraging their vast pre-existing knowledge and reasoning capabilities. One direction explores LLMs as end-to-end solvers by fine-tuning them on large-scale datasets of problems paired with their solutions [1, 2]. Another approach employs LLMs as automated problem modelers that translate natural language into executable code for traditional solvers [5, 3]. A third direction positions the LLM as a hyper-heuristic generator that iteratively reflects upon and improves heuristic rule-code [21]. Despite their potential, early works reveal critical challenges: the risk of generating infeasible solutions (“hallucinations”), high computational costs, and a primary focus on static problems, leaving their application to complex, dynamic scheduling environments largely unexploredc.
@@
Hierarchical Memory and Reasoning Paradigms: A significant body of work explores hierarchical paradigms to structure and access information beyond the fixed context window. A prominent direction involves creating hierarchical memory systems, such as MemGPT’s OS-inspired information paging [41] or the semantic-level memory structures in MemoryBank and H-MEM that enable targeted retrieval [70, 50]. A related approach, seen in HIAGENT, employs hierarchy for task decomposition to manage an agent’s working memory during long-horizon tasks [20]. Diverging from these methods, which primarily focus on organizing and retrieving an agent’s past, our proposed ReflecSched framework introduces a hierarchical reflection mechanism that functions as a dynamic, planning-time operator. Instead of retrieving historical data, it prospectively explores the future by leveraging multi-level, heuristic-driven simulations to analyze potential decision pathways. The hierarchy in ReflecSched is thus one of planning abstraction distilling numerical simulations into a transient, textual Strategic Experience to guide the next action, rather than creating a persistent memory store.
@@ 3.1 Dynamic Flexible Job-Shop Scheduling
DFJSP is the problem of assigning operations from multiple jobs to a set of machines, subject to real-time stochastic events. It is characterized by two features: flexibility, which allows each operation to be processed on a subset of candidate machines, and dynamism, which refers to the occurrence of real-time stochastic events such as new job arrivals or machine breakdowns. The objective is to generate a valid schedule that minimizes the makespan, defined as the maximum completion time among all jobs [12]. Figure 1 illustrates an example of such a schedule, visualized as a Gantt chart. A valid schedule must satisfy both precedence constraints between operations of the same job and the resource constraints of the machines.
In this work, we model a rich set of dynamic events to capture the stochastic nature of modern manufacturing environments. These include new job arrivals, machine breakdowns and subsequent repairs, job cancellations, and the introduction of high-priority emergency jobs. This set of events is representative of the primary disruptions studied in the DFJSP literature [8, 40, 53]. More importantly, the core architecture of ReflecSched is designed to be event-agnostic and generalizable. Its strategic Hierarchical Reflection module is invoked by an event-driven trigger, which responds to any significant state change rather than to a specific event type. Consequently, the framework can be readily extended to handle other stochastic events, such as variable processing times [46] or rush order arrivals [43], without requiring modifications to its fundamental logic.
Table 1: Notation for the DFJSP Mathematical Formulation
Symbol Description
Sets and Indices
$t$ Discrete decision step index, $t=0,1,2,$...
$\tau$ Continuous physical time index, $\tau\in[0,\infty)$.
$\mathcal{J}$ Set of active jobs currently in the system, $\{1,...,n\}$.
$\mathcal{M}$ Set of available machines, $\{1,...,m\}$.
$\mathcal{O}$ Set of all operations to be scheduled.
$\mathcal{O}_{i}$ Set of operations for job $J_{i}\in\mathcal{J}$.
$\mathcal{M}_{ij}$ Set of candidate machines for operation $O_{ij}$.
Parameters and Constants
$n_{i}$ Number of operations for job $J_{i}$.
$p_{ijk}$ Processing time of $O_{ij}$ on machine $k\in\mathcal{M}_{ij}$.
$a_{i}$ Arrival time of job $J_{i}$.
$M_{big}$ A sufficiently large positive number (a “Big-M” constant).
Decision Variables
$x_{ijk}$ Binary: 1 if $O_{ij}$ is assigned to machine $k$.
Auxiliary and State Variables
$s_{ij}$ Start time of operation $O_{ij}$.
$c_{ij}$ Completion time of operation $O_{ij}$.
$r_{k}$ Earliest time machine $k$ becomes free.
$C_{max}$ The makespan, defined as $\max_{J_{i}\in\mathcal{J}}\{c_{i,n_{i}}\}$.
@@ 3.2 Mathematical Formulation of the DFJSP
3.2.1 Model Assumptions
To ensure clarity and specify the applicable scope of the proposed mathematical model, we explicitly state the following assumptions regarding the DFJSP environment [68]. These assumptions define the boundaries of resource constraints, temporal dynamics, and event handling logic:
To define the operational scope of the proposed mathematical model, we stipulate the following assumptions regarding the DFJSP environment [69]. These assumptions delineate the parameters of resource constraints, temporal dynamics, and the logic governing dynamic events:
Resource and Task Constraints: Each job follows a predefined sequence of operations, with each operation occupying a single machine at any given time. Machines constitute the primary capacity constraints and are assumed to have infinite buffer capacity. Auxiliary resources, such as operators and tools, are considered non-constraining and remain readily available.
Temporal Specifications: Setup times are assumed to be sequence-independent and are incorporated into the operation processing times. Similarly, inter-machine transfer times are considered negligible or are subsumed under the processing durations. All processing times are deterministic and remain invariant throughout the execution of an operation.
Dynamic Event Logic: The scheduler operates in an online environment where the timing and characteristics of stochastic events, such as job arrivals and machine breakdowns, are unknown a priori. Operations are non-preemptive except when interrupted by equipment failures. In such instances, we assume a preempt-resume logic: the accumulated processing progress is preserved, and the operation resumes as soon as the machine becomes available again.
@@ 3.2.2 Objective Function and Constraints
$s_{i^{\prime}j^{\prime}}$ $\geq c_{ij}-{\color[rgb]{0,0,1}M_{big}}\cdot(3-x_{ijk}-x_{i^{\prime}j^{\prime}k}-y_{ij,i^{\prime}j^{\prime}k})$ (7)
$s_{ij}$ $\geq c_{i^{\prime}j^{\prime}}-{\color[rgb]{0,0,1}M_{big}}\cdot(2+y_{ij,i^{\prime}j^{\prime}k}-x_{ijk}-x_{i^{\prime}j^{\prime}k})$ (8)
@@ 3.2.3 Formalization of Dynamic Events
The state of the system at time $\tau$ is defined by the tuple $S_{\tau}=(\mathcal{J}_{\tau},\{a_{i}\},\{r_{k}\},\mathcal{O}_{int})$. A dynamic event occurring at time $\tau$ triggers an instantaneous state transition from $S_{\tau}$ to $S_{\tau^{+}}$, where $\tau^{+}$ denotes the time immediately following the event execution. The specific update logic for each event type is formalized below [7].
Job Arrival: At physical time $\tau=\tau_{new}$, a new job $J_{new}$ with its own set of operations $\mathcal{O}_{new}$ is introduced into the system [47]. The state is updated instantaneously as follows:
$\mathcal{J}_{\tau^{+}}$ $\leftarrow\mathcal{J}_{\tau}\cup\{J_{new}\}$ (9)
$\mathcal{O}_{\tau^{+}}$ $\leftarrow O_{\tau}\cup O_{new}$ (10)
The emergency status of a job is a characteristic handled by the scheduling policy via prioritization rather than being a direct modification to the model’s constraints.
Job Cancellation: At physical time $\tau=\tau_{cancel}$, a job $J_{cancel}$ is removed from the system. This action removes the job from the set of active jobs $\mathcal{J}$ and all of its corresponding uncompleted operations from the set of operations to be scheduled $\mathcal{O}_{cancel}$. The historical record of any completed operations of $J_{cancel}$ is preserved, but they are no longer relevant for future scheduling decisions.
@@
$\mathcal{J}_{\tau^{+}}$ $\leftarrow\mathcal{J}_{\tau}\setminus\{J_{cancel}\}$ (11)
$O_{\tau^{+}}$ $\leftarrow O_{\tau}\setminus O_{cancel}$ (12)
If an operation from $\mathcal{O}_{cancel}$ was in process on a machine $k$ at $\tau$, it is preempted. The machine is immediately freed, and its availability time $r_{k}$ is updated to $\tau$.
Machine Breakdown: At physical time $\tau=\tau_{break}$, a machine $k\in\mathcal{M}$ fails and becomes unavailable [30]. The immediate consequence is that the machine’s earliest availability time is postponed until its projected repair time, $\tau_{repair}$. This is formalized as:
$r_{k}=\tau_{repair}$ (13)
This update affects all subsequent scheduling decisions, as no new operations can be assigned to machine $k$ until $\tau_{repair}$.
If an operation $O_{ij}$ was being processed on machine $k$ at this time i.e., $s_{ij}<\tau_{break}<c_{ij}$, it is immediately interrupted. The operation is formally moved to a set of interrupted operations, $\mathcal{O}_{int}$, to signify its special status. It retains its assignment to machine $k$ and awaits the machine’s repair to resume processing.
Machine Repair: At physical time $\tau=\tau_{repair}$, machine $k$ becomes operational again. Any operation $O_{ij}\in\mathcal{O}_{int}$ that was interrupted on this machine can now resume. Upon repair, $O_{ij}$ is removed from the set $\mathcal{O}_{int}$.
@@
$c^{\prime}_{ij}=\tau+p_{ijk}-(\tau_{break}-s_{ij})$ (14)
where $\tau_{break}-s_{ij}$ represents the processing work completed before the interruption. This new completion time $c^{\prime}_{ij}$ subsequently serves as the basis for re-evaluating the precedence constraints for all downstream operations of job $J_{i}$. The machine $k$ is also now available for scheduling other operations from time $\tau_{repair}$ onwards.
3.3 Industrial Application of Semiconductor Cluster Tool Scheduling
The DFJSP framework investigated in this study is motivated by the operational management of cluster tools within semiconductor wafer fabrication facilities. As illustrated in Figure 2, a cluster tool is a representative manufacturing unit comprising several heterogeneous processing chambers, denoted as PM1 through PM5, which surround a central multi-link handling robot [42]. Wafers enter the system through load lock interfaces, after which the robot executes the sequence of transport steps required to move them between specific modules. These chambers perform high-precision processes, such as chemical vapor deposition or reactive ion etching, where scheduling complexity is dictated by a set of technological constraints that transcend traditional job-shop models.
Machine eligibility is determined by wafer recipes, as only a specific subset of process modules is configured with the gas chemistry, RF power, or thermal settings required for a given technological step [58]. Furthermore, re-entrant flow patterns are prevalent in this environment; wafers must frequently revisit modules such as PM2 or PM3 to undergo multi-layer processing at various stages of the production cycle. Yield management introduces a third critical constraint: residency time limits. These constraints mandate that the handling robot extracts wafers from a chamber immediately upon processing completion, as prolonged exposure to the chamber environment renders the wafer surface vulnerable to oxidation or chemical contamination. Consequently, the scheduler must precisely synchronize robot transport with chamber cycle times to mitigate defects induced by processing delays.
The operational dynamism of this fabrication environment is further characterized by frequent perturbations that necessitate the real-time reconfiguration of pre-planned schedules. These disturbances encompass the stochastic arrival of high-priority “hot lots” requiring immediate prioritization to satisfy stringent delivery constraints, equipment failures resulting from parameter drifts, and job cancellations triggered by upstream quality excursions. When a processing module becomes unavailable, the system must autonomously identify alternative routings through compatible modules to ensure production continuity. Addressing these challenges requires a scheduling framework capable of non-myopic reasoning to optimize throughput while adhering to the physical constraints of the equipment. Consequently, the coordination of semiconductor cluster tools provides a compelling case for the proposed hierarchical reflection mechanism, as it offers a structured paradigm for balancing transient efficiency with long-term schedule robustness [4].
Figure 2: Architectural schematic of a semiconductor cluster tool utilized in the JMS-Bench dataset. The system integrates five heterogeneous process modules (PM1 to PM5) with a central transfer module and dual-arm robot. Wafers are loaded from cassettes and navigate through multi-stage recipes involving re-entrant flows and strict residency time constraints.